please help me with my homework: using exact values, verify that cos 60 degrees = 1 - 2 sin^2 30 degrees.
Answer by stanbon(48558)

verify that cos 60 degrees = 1 - 2 sin^2 30 degrees.
---------
cos 60 = 1/2
sin 30 = 1/2
Your Problem:
1/2 = 1 - 2(1/2)^2
1/2 = 1 - 2*(1/4)
1/2 = 1 - 1/2
1/2 = 1/2
==============
Cheers,
Stan H.